OpenStudy (anonymous):

The formula d = rt gives the distance traveled in time t at rate r. A delivery truck drives for 1.5 hours and travels 76 miles. What was the average speed of the truck?

jhonyy9 (jhonyy9):

so there is formula d=r*t where d is distance ,time t and rate r what you have got in your exercise time t = 1,5 hours and travels 76 miles = distance d what is the average speed of the truck r d=r*t so from this you get r=d/t - before you make the calculi i think you need convert this 1,5 hours in minutes - so you know that 1 hour is = 60 minutes than 1,5 hours will be = 90 minutes - than using this formula you will get r=76/90 = 0,84 miles per minutes - but you need giving it in miles/hours so than you need just multiply 0,84 by 60 what is = 50,4 miles - so what mean that the average speed of the truck r = 50,4 miles/hours - hope so much that this will help you and that this is understandably sure right easy